[Brahms-l] Shift report 20060612 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Mon Jun 12 2006 - 16:26:06 EDT
Shifter leader was kh
Shifters were nk


Beam on and off during most of the day always ending with unexpected beam aborts.  Total usable beam time was about 3 hrs
Took data during the time on with FS 3 deg 1/2A and MRS 45 deg 6kG (1050) B
Angelika did a vernier scan which we got 2.75 out of the 4 scans in run 16099, a min bias run.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Jun 12 16:29:05 2006

This archive was generated by hypermail 2.1.8 : Mon Jun 12 2006 - 16:29:19 EDT